OpenStudy (kise_ryouta):

What is the value of g(−3) when g(x)=2x−2 ? Enter your answer in the box. g(−3)=

OpenStudy (kise_ryouta):

i think its -6

OpenStudy (tkhunny):

You "think"? That;s not good enough. Please DEMONSTRATE it. g(-3) = 2(-3) - 2 = ??

OpenStudy (kise_ryouta):

wait bc 2*-3 is -6 but -2 it would be -4 right

OpenStudy (tkhunny):

Why would negative 2 increase the value? Make sure you know which way things are going.

OpenStudy (kise_ryouta):

no im saying subtract 2

OpenStudy (kise_ryouta):

so it would be -8

OpenStudy (kise_ryouta):

i think idk im confused

OpenStudy (tkhunny):

2(-3) - 2 = -6 - 2 = -8 See, you have it. :-)

OpenStudy (kise_ryouta):

lol okay i was overthinking it a little but can you help me with more

OpenStudy (tkhunny):

Please make a new thread. Someone will help you if you show your work up front. Don't just post the problem.